# Deep End-to-end Adaptive k-Space Sampling, Reconstruction, and Registration for Dynamic MRI

### v1

@@ 0.C.1 Additional Tables
$\alpha$ $\beta$ SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
vSHARP Adaptive Learned 1 1 0.859 $\pm$ 0.082 31.96 $\pm$ 3.65 0.059 $\pm$ 0.042 0.859 $\pm$ 0.082 32.04 $\pm$ 3.67 0.058 $\pm$ 0.041 0.857 $\pm$ 0.082 32.07 $\pm$ 3.68 0.058 $\pm$ 0.041
vSHARP Adaptive Learned (L) 1 1 0.860 $\pm$ 0.081 31.93 $\pm$ 3.57 0.059 $\pm$ 0.040 0.860 $\pm$ 0.080 32.03 $\pm$ 3.58 0.058 $\pm$ 0.039 0.858 $\pm$ 0.079 32.03 $\pm$ 3.57 0.057 $\pm$ 0.039
vSHARP Adaptive Voxelmorph 1 1 0.839 $\pm$ 0.085 31.10 $\pm$ 3.40 0.069 $\pm$ 0.044 0.835 $\pm$ 0.086 31.05 $\pm$ 3.42 0.070 $\pm$ 0.044 0.830 $\pm$ 0.086 30.94 $\pm$ 3.42 0.071 $\pm$ 0.044
vSHARP Adaptive Transfmorph 1 1 0.835 $\pm$ 0.092 30.46 $\pm$ 3.48 0.083 $\pm$ 0.060 0.834 $\pm$ 0.091 30.53 $\pm$ 3.48 0.082 $\pm$ 0.059 0.835 $\pm$ 0.091 30.65 $\pm$ 3.51 0.080 $\pm$ 0.058
vSHARP Adaptive Learned 1 3 0.891 $\pm$ 0.067 34.36 $\pm$ 3.82 0.034 $\pm$ 0.023 0.887 $\pm$ 0.069 34.18 $\pm$ 3.83 0.035 $\pm$ 0.023 0.882 $\pm$ 0.070 33.93 $\pm$ 3.80 0.037 $\pm$ 0.024
vSHARP Adaptive Learned 1 1 0.860 $\pm$ 0.080 32.00 $\pm$ 3.58 0.058 $\pm$ 0.041 0.859 $\pm$ 0.080 32.05 $\pm$ 3.59 0.058 $\pm$ 0.040 0.857 $\pm$ 0.080 32.04 $\pm$ 3.59 0.058 $\pm$ 0.039
vSHARP Adaptive Learned 1 1 0.858 $\pm$ 0.083 31.87 $\pm$ 3.61 0.060 $\pm$ 0.041 0.858 $\pm$ 0.082 31.95 $\pm$ 3.62 0.059 $\pm$ 0.040 0.855 $\pm$ 0.082 31.90 $\pm$ 3.58 0.059 $\pm$ 0.040
vSHARP Adaptive Learned 1 1 0.855 $\pm$ 0.083 31.56 $\pm$ 3.59 0.065 $\pm$ 0.046 0.855 $\pm$ 0.082 31.65 $\pm$ 3.59 0.064 $\pm$ 0.045 0.852 $\pm$ 0.082 31.64 $\pm$ 3.56 0.063 $\pm$ 0.044
vSHARP Equi NA Learned 1 1 0.849 $\pm$ 0.086 31.13 $\pm$ 3.53 0.071 $\pm$ 0.049 0.851 $\pm$ 0.086 31.25 $\pm$ 3.53 0.069 $\pm$ 0.048 0.847 $\pm$ 0.087 31.24 $\pm$ 3.52 0.069 $\pm$ 0.047
vSHARP Equi NA Learned 1 1 0.849 $\pm$ 0.087 31.12 $\pm$ 3.55 0.072 $\pm$ 0.050 0.849 $\pm$ 0.087 31.18 $\pm$ 3.52 0.070 $\pm$ 0.049 0.844 $\pm$ 0.087 31.15 $\pm$ 3.49 0.070 $\pm$ 0.048
vSHARP $k$t-Equi NA Learned 1 1 0.847 $\pm$ 0.086 31.19 $\pm$ 3.56 0.070 $\pm$ 0.049 0.847 $\pm$ 0.085 31.27 $\pm$ 3.53 0.069 $\pm$ 0.047 0.839 $\pm$ 0.087 31.20 $\pm$ 3.55 0.069 $\pm$ 0.047
vSHARP $k$t-Equi NA Learned 1 1 0.847 $\pm$ 0.087 31.03 $\pm$ 3.54 0.073 $\pm$ 0.050 0.848 $\pm$ 0.086 31.11 $\pm$ 3.50 0.071 $\pm$ 0.048 0.841 $\pm$ 0.089 31.06 $\pm$ 3.51 0.071 $\pm$ 0.048
vSHARP Optimized Learned 1 1 0.863 $\pm$ 0.081 32.20 $\pm$ 3.60 0.055 $\pm$ 0.037 0.863 $\pm$ 0.080 32.26 $\pm$ 3.58 0.054 $\pm$ 0.036 0.861 $\pm$ 0.080 32.25 $\pm$ 3.56 0.054 $\pm$ 0.035
VarNet Adaptive Learned 1 1 0.835 $\pm$ 0.083 31.42 $\pm$ 3.25 0.060 $\pm$ 0.031 0.811 $\pm$ 0.087 30.57 $\pm$ 3.07 0.071 $\pm$ 0.031 0.807 $\pm$ 0.089 30.34 $\pm$ 2.99 0.074 $\pm$ 0.031
vSHARP Adaptive OFILK NA 1 NA 0.829 $\pm$ 0.089 30.41 $\pm$ 3.25 0.081 $\pm$ 0.053 0.827 $\pm$ 0.089 30.42 $\pm$ 3.23 0.081 $\pm$ 0.052 0.825 $\pm$ 0.089 30.41 $\pm$ 3.20 0.081 $\pm$ 0.051
vSHARP Adaptive OFTVL1 NA 1 NA 0.817 $\pm$ 0.071 28.11 $\pm$ 2.32 0.129 $\pm$ 0.064 0.814 $\pm$ 0.071 28.02 $\pm$ 2.28 0.132 $\pm$ 0.065 0.811 $\pm$ 0.070 27.93 $\pm$ 2.27 0.134 $\pm$ 0.065
vSHARP Adaptive DEMONS NA 1 NA 0.848 $\pm$ 0.087 30.58 $\pm$ 3.74 0.085 $\pm$ 0.072 0.847 $\pm$ 0.087 30.57 $\pm$ 3.72 0.085 $\pm$ 0.070 0.844 $\pm$ 0.087 30.55 $\pm$ 3.71 0.085 $\pm$ 0.069
vSHARP Equi NA OFILK NA 1 NA 0.827 $\pm$ 0.091 30.37 $\pm$ 3.29 0.082 $\pm$ 0.054 0.826 $\pm$ 0.091 30.40 $\pm$ 3.26 0.081 $\pm$ 0.052 0.821 $\pm$ 0.092 30.33 $\pm$ 3.21 0.082 $\pm$ 0.050
vSHARP Equi NA OFTVL1 NA 1 NA 0.819 $\pm$ 0.072 28.12 $\pm$ 2.32 0.128 $\pm$ 0.063 0.817 $\pm$ 0.072 28.05 $\pm$ 2.29 0.130 $\pm$ 0.063 0.808 $\pm$ 0.074 27.81 $\pm$ 2.27 0.136 $\pm$ 0.064
vSHARP Equi NA DEMONS NA 1 NA 0.848 $\pm$ 0.088 30.54 $\pm$ 3.76 0.086 $\pm$ 0.074 0.847 $\pm$ 0.089 30.55 $\pm$ 3.72 0.085 $\pm$ 0.071 0.841 $\pm$ 0.090 30.49 $\pm$ 3.65 0.085 $\pm$ 0.067
vSHARP $k$t-Equi NA OFILK NA 1 NA 0.826 $\pm$ 0.091 30.36 $\pm$ 3.30 0.082 $\pm$ 0.054 0.824 $\pm$ 0.091 30.37 $\pm$ 3.28 0.082 $\pm$ 0.053 0.816 $\pm$ 0.093 30.30 $\pm$ 3.24 0.082 $\pm$ 0.051
vSHARP $k$t-Equi NA OFTVL1 NA 1 NA 0.815 $\pm$ 0.073 28.11 $\pm$ 2.35 0.129 $\pm$ 0.064 0.810 $\pm$ 0.073 27.98 $\pm$ 2.33 0.133 $\pm$ 0.065 0.798 $\pm$ 0.076 27.67 $\pm$ 2.32 0.142 $\pm$ 0.067
vSHARP $k$t-Equi NA DEMONS NA 1 NA 0.847 $\pm$ 0.088 30.54 $\pm$ 3.77 0.086 $\pm$ 0.073 0.845 $\pm$ 0.088 30.53 $\pm$ 3.73 0.086 $\pm$ 0.071 0.837 $\pm$ 0.090 30.48 $\pm$ 3.68 0.085 $\pm$ 0.065
Table 3: Quantitative results for various configurations on the aorta inference set (under phase-specific settings).
Reconstruction Model Sampling Type Sampling Initialization Registration Module Joint Loss Loss Weights 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 ×
$\alpha$ $\beta$ SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
vSHARP Adaptive Learned 1 1 0.837 $\pm$ 0.088 30.79 $\pm$ 3.45 0.075 $\pm$ 0.051 0.825 $\pm$ 0.090 30.49 $\pm$ 3.33 0.079 $\pm$ 0.049 0.809 $\pm$ 0.092 30.04 $\pm$ 3.23 0.085 $\pm$ 0.049
vSHARP Adaptive Learned 1 1 0.837 $\pm$ 0.086 30.85 $\pm$ 3.38 0.074 $\pm$ 0.049 0.827 $\pm$ 0.086 30.56 $\pm$ 3.29 0.077 $\pm$ 0.047 0.813 $\pm$ 0.088 30.18 $\pm$ 3.17 0.082 $\pm$ 0.046
vSHARP Adaptive Learned 1 1 0.841 $\pm$ 0.087 30.91 $\pm$ 3.49 0.074 $\pm$ 0.052 0.832 $\pm$ 0.086 30.77 $\pm$ 3.41 0.075 $\pm$ 0.050 0.820 $\pm$ 0.087 30.43 $\pm$ 3.30 0.079 $\pm$ 0.049
vSHARP Adaptive Learned 1 1 0.840 $\pm$ 0.087 30.88 $\pm$ 3.43 0.074 $\pm$ 0.049 0.830 $\pm$ 0.087 30.67 $\pm$ 3.32 0.076 $\pm$ 0.048 0.816 $\pm$ 0.089 30.29 $\pm$ 3.23 0.080 $\pm$ 0.047
vSHARP Equi NA Learned 1 1 0.837 $\pm$ 0.090 30.67 $\pm$ 3.50 0.078 $\pm$ 0.054 0.828 $\pm$ 0.091 30.59 $\pm$ 3.40 0.079 $\pm$ 0.052 0.803 $\pm$ 0.093 29.90 $\pm$ 3.14 0.088 $\pm$ 0.051
vSHARP Equi NA Learned 1 1 0.840 $\pm$ 0.090 30.83 $\pm$ 3.55 0.076 $\pm$ 0.054 0.831 $\pm$ 0.091 30.73 $\pm$ 3.46 0.076 $\pm$ 0.052 0.805 $\pm$ 0.094 30.07 $\pm$ 3.24 0.085 $\pm$ 0.051
VarNet Adaptive Learned 1 1 0.782 $\pm$ 0.089 29.56 $\pm$ 2.80 0.086 $\pm$ 0.028 0.769 $\pm$ 0.090 29.15 $\pm$ 2.69 0.094 $\pm$ 0.029 0.763 $\pm$ 0.089 28.97 $\pm$ 2.66 0.098 $\pm$ 0.030
vSHARP Adaptive OFILK NA 1 NA 0.810 $\pm$ 0.092 29.95 $\pm$ 3.09 0.087 $\pm$ 0.052 0.795 $\pm$ 0.092 29.53 $\pm$ 2.93 0.094 $\pm$ 0.050 0.776 $\pm$ 0.092 29.04 $\pm$ 2.80 0.102 $\pm$ 0.048
vSHARP Adaptive OFTVL1 NA 1 NA 0.792 $\pm$ 0.072 27.52 $\pm$ 2.20 0.145 $\pm$ 0.065 0.771 $\pm$ 0.072 26.97 $\pm$ 2.12 0.162 $\pm$ 0.069 0.747 $\pm$ 0.071 26.44 $\pm$ 2.07 0.181 $\pm$ 0.073
vSHARP Adaptive DEMONS NA 1 NA 0.832 $\pm$ 0.090 30.13 $\pm$ 3.57 0.091 $\pm$ 0.071 0.817 $\pm$ 0.091 29.72 $\pm$ 3.39 0.096 $\pm$ 0.069 0.800 $\pm$ 0.093 29.28 $\pm$ 3.23 0.103 $\pm$ 0.066
vSHARP Equi NA OFILK NA 1 NA 0.820 $\pm$ 0.091 30.22 $\pm$ 3.27 0.084 $\pm$ 0.054 0.806 $\pm$ 0.092 29.98 $\pm$ 3.14 0.088 $\pm$ 0.053 0.778 $\pm$ 0.093 29.38 $\pm$ 2.92 0.097 $\pm$ 0.050
vSHARP Equi NA OFTVL1 NA 1 NA 0.809 $\pm$ 0.073 28.00 $\pm$ 2.33 0.132 $\pm$ 0.064 0.792 $\pm$ 0.074 27.64 $\pm$ 2.27 0.141 $\pm$ 0.066 0.758 $\pm$ 0.074 26.87 $\pm$ 2.17 0.165 $\pm$ 0.068
vSHARP Equi NA DEMONS NA 1 NA 0.840 $\pm$ 0.090 30.40 $\pm$ 3.73 0.088 $\pm$ 0.074 0.826 $\pm$ 0.092 30.14 $\pm$ 3.62 0.091 $\pm$ 0.072 0.797 $\pm$ 0.094 29.50 $\pm$ 3.34 0.100 $\pm$ 0.067
vSHARP Optimized Learned 1 1 0.846 $\pm$ 0.085 31.15 $\pm$ 3.51 0.070 $\pm$ 0.049 0.838 $\pm$ 0.086 30.85 $\pm$ 3.41 0.073 $\pm$ 0.048 0.828 $\pm$ 0.086 30.59 $\pm$ 3.31 0.076 $\pm$ 0.047

### v2

@@ C.2 Additional Tables
$\mathcal{L}_{\mathrm{sim}},\mathcal{L}_{\mathrm{reg}}$ $\alpha$ $\beta$ $\mathcal{L}_{\mathrm{smooth}}$ E2E Loss SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 $0.877\pm 0.050$ $31.68\pm 2.86$ $0.037\pm 0.019$ $0.867\pm 0.052$ $31.41\pm 2.76$ $0.038\pm 0.019$ $0.854\pm 0.054$ $31.01\pm 2.69$ $0.042\pm 0.020$ 10.06
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 $0.882\pm 0.049$ $31.90\pm 2.92$ $0.035\pm 0.019$ $0.874\pm 0.051$ $31.75\pm 2.84$ $0.036\pm 0.019$ $0.862\pm 0.054$ $31.40\pm 2.75$ $0.039\pm 0.020$ 10.08
Learned vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 $0.879\pm 0.050$ $31.62\pm 2.98$ $0.038\pm 0.021$ $0.868\pm 0.053$ $31.51\pm 2.88$ $0.038\pm 0.020$ $0.848\pm 0.057$ $30.93\pm 2.72$ $0.043\pm 0.021$ 9.96
Learned VarNet Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 $0.837\pm 0.056$ $30.92\pm 2.40$ $0.041\pm 0.015$ $0.832\pm 0.057$ $30.77\pm 2.40$ $0.043\pm 0.016$ $0.829\pm 0.058$ $30.66\pm 2.38$ $0.044\pm 0.016$ 10.54
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.868\pm 0.049$ $31.53\pm 2.68$ $0.037\pm 0.019$ $0.856\pm 0.052$ $31.15\pm 2.65$ $0.041\pm 0.019$ $0.842\pm 0.056$ $30.70\pm 2.62$ $0.045\pm 0.021$ 24.93
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.873\pm 0.045$ $30.01\pm 2.78$ $0.055\pm 0.031$ $0.861\pm 0.049$ $29.55\pm 2.76$ $0.061\pm 0.034$ $0.848\pm 0.053$ $29.01\pm 2.77$ $0.069\pm 0.038$ 29.98
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.873\pm 0.050$ $30.36\pm 2.80$ $0.049\pm 0.024$ $0.863\pm 0.053$ $30.19\pm 2.78$ $0.051\pm 0.025$ $0.852\pm 0.056$ $29.93\pm 2.73$ $0.054\pm 0.026$ 29.11
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.874\pm 0.047$ $31.69\pm 2.69$ $0.036\pm 0.018$ $0.862\pm 0.050$ $31.34\pm 2.61$ $0.039\pm 0.019$ $0.849\pm 0.053$ $30.94\pm 2.57$ $0.042\pm 0.020$ 25.51
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.877\pm 0.044$ $30.18\pm 2.86$ $0.053\pm 0.030$ $0.867\pm 0.047$ $29.76\pm 2.73$ $0.058\pm 0.033$ $0.855\pm 0.050$ $29.31\pm 2.70$ $0.064\pm 0.035$ 29.90
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.877\pm 0.048$ $30.50\pm 2.79$ $0.048\pm 0.024$ $0.868\pm 0.052$ $30.34\pm 2.76$ $0.049\pm 0.024$ $0.857\pm 0.054$ $30.14\pm 2.71$ $0.051\pm 0.025$ 28.91
DEMONS vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.877\pm 0.046$ $31.82\pm 2.74$ $0.035\pm 0.018$ $0.861\pm 0.050$ $31.43\pm 2.65$ $0.038\pm 0.018$ $0.838\pm 0.055$ $30.72\pm 2.50$ $0.044\pm 0.019$ 26.66
DEMONS vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.882\pm 0.044$ $30.40\pm 2.99$ $0.051\pm 0.031$ $0.868\pm 0.047$ $30.06\pm 2.87$ $0.054\pm 0.031$ $0.847\pm 0.052$ $29.32\pm 2.72$ $0.064\pm 0.035$ 30.04
DEMONS vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 NA $0.880\pm 0.048$ $30.55\pm 2.83$ $0.047\pm 0.024$ $0.865\pm 0.052$ $30.32\pm 2.77$ $0.050\pm 0.024$ $0.845\pm 0.056$ $29.89\pm 2.64$ $0.054\pm 0.025$ 29.29
Learned vSHARP Optimized $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1 1 $0.886\pm 0.047$ $32.17\pm 2.87$ $0.033\pm 0.018$ $0.879\pm 0.048$ $31.96\pm 2.80$ $0.034\pm 0.018$ $0.870\pm 0.050$ $31.71\pm 2.72$ $0.036\pm 0.018$ 10.83